OpenStudy (anonymous):

solve this proportion: 117/x=11/44

OpenStudy (anonymous):

look at it as 11x = 44(117), from here it is easy to solve for x,try it

OpenStudy (anonymous):

A more clear explanation to what @amiliver did is cross multiply, do you know how to do that?

OpenStudy (anonymous):

our purpose here is again to solve for "x"

OpenStudy (anonymous):

To set it our clearly, I will show you. \[\frac{117}{x}=\frac{11}{44}\] Hence the word "Cross" in cross multiply, we multiply diagonally. (11 times x) = (117 times 44) 11 x =5148 x = 5148/11

OpenStudy (anonymous):

You are only allowed to cross multiply when there is one fraction equal to another fraction just like this example.

OpenStudy (anonymous):

ok thanks for the tip

OpenStudy (anonymous):

no problem, I am the master of simple maths lol, I'm joking.